OpenStudy (anonymous):

how to use the nth root theorem to find the three roots of z^3 = -343

OpenStudy (solomonzelman):

\[(-7)^3=-343\]

OpenStudy (anonymous):

there are 3 roots

OpenStudy (solomonzelman):

WHAT?!?! No.... I just gave you a hint.

OpenStudy (anonymous):

factor and use the quadratic formula for the two that are not obvious

OpenStudy (anonymous):

No I need use the nth theorem to find them.

OpenStudy (anonymous):

what on earth is an "nth root theorem"?

OpenStudy (anonymous):

demoivre's theorem

OpenStudy (anonymous):

k divide the unit circle in to three equal parts, with one of the parts at \(-1\) find the complex numbers multiply the result by \(7\)

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay I get that but how do I turn -343 into an angle?

OpenStudy (anonymous):

you do not it is a number, not an angle

OpenStudy (anonymous):

finding the cube roots of \(-343\) is identical to finding the cube roots of \(-1\) except you multiply the results by \(7\)

OpenStudy (anonymous):

if you want the angle for \(-1\) look at the unit circle in the complex plane. it should be pretty clear that the angle is question is \(\pi\) or if you are working in degrees, \(180\)

OpenStudy (anonymous):

I don't really understand. In my book they show how to do it in polar form but I don't know how to get -343 into polar . Do you know how to convert it because I think that would help.

OpenStudy (anonymous):

\[-343=-343+0i\] The real and imaginary parts, in polar form: \[\begin{cases}-343=r\cos\theta\\ 0=r\sin\theta\end{cases}~~\Rightarrow~~r=343,~\theta=\pi\] since -343 lies on the real axis. \(\theta=\pi\) because \(r\) can't be negative. So, \[\large z^3=343e^{i\pi}~~\iff~~z=343^{1/3}e^{i\pi/3}\] Of course, you also have to account for the complex roots, which have the form \[{\huge 343^{1/3}e^{i\left(\frac{\pi}{3}+\frac{2\pi}{3}\right)}}~~\text{and}~~{\huge 343^{1/3}e^{i\left(\frac{\pi}{3}+\frac{4\pi}{3}\right)}}\]